Presidio, Inc. produces one model of mountain bike. Partial information for the company ollows Required: 1. Complete Presidios cost data table. (Round your Cost per Unit answers to 2 decim places.) Bikes Produced and Sold 450 Units 910 Units Total costs 16a Units Variable costs S 144,000 Fixed costs per year Total costs 144.000 Cost per unit Variable cost per unit Fixed cost per unit Total cost per unit 0.00 S 53800 S 0.00 2. sales level indicated in the cost data table assuming the company sells each bike for (Round your Margin Ratio percentage answers to 2 decimal places (le. .1234 should be entered as 12.34%.)) Calculate Presidios contribution margin ratio and its total contribution margin at each $640 450 Units 910 Units 1628 Units Total Contribution Margin Contribution Margin Ratio 3. Calculate net operating income (loss) at each of the sales levels assuming a sales price of Shan (Round vour answers to the nearest whole dollaramount.)

media%2F199%2F199ab432-2ab3-4cbb-87cb-77
Presidio, Inc. produces one model of mountain bike. Partial information for the company follows:

Required:
1. Complete Presidio’s cost data table.(Round your Cost per Unit answers to 2 decimal places.)



2. Calculate Presidio’s contribution margin ratio and its total contribution margin at each sales level indicated in the cost data table assuming the company sells each bike for $640. (Round your Margin Ratio percentage answers to 2 decimal places (i.e. .1234 should be entered as 12.34%.))



3. Calculate net operating income (loss) at each of the sales levels assuming a sales price of $640. (Round your answers to the nearest whole dollar amount.
Answer #1

(1) Cost Data Table :-

450 Units

910 Units

1628 Unts

Total Costs

Variable cost

144000

(910 * 320) = 291200

(1628 * 320) = 520960

Fixed cost per year

198380**

198380**

198380**

Total Costs

342380

489580

719340

Cost per unit

Variable cost pu

320*

320*

320*

Fixed cost pu

440.84

218

121.86

Total cost pu

760.84

538

441.86

*Variable cost per unit = $144000/450 units = $320

VC per unit will remain same

**Fixed Cost :-

In 910 Units:-

Cost pu = $538

Total Cost = $538 * 910 units = $489580

Variable cost for 910 units = 910 units * $320 = $291200

Fixed Cost = Total cost – Variable cost

       = 489580 – 291200 = 198380

Fixed cost will remain same in all the above cases irrespective of the no of units produced & sell

(2) Contribution Margin Ratio & its Total contribution Margin :-

450 Units

910 Units

1628 Units

Total contribution Margin

144000

291200

520960

Contribution Margin Ratio

50%

50%

50%

SP pu = $640

Contribution Margin = Sale Value – Variable cost

Contribution Margin Ratio = (Contribution Margin / Sale Value) * 100

450 Units

910 Units

1628 Units

Sale value                                 (A)

(450 * $640)

=288000

(910 * $640)

=582400

(1628 * $640)

=1041920

(-) Variable Cost

144000

291200

520960

Contribution Margin              (B)

144000

291200

520960

Contribution Margin Ratio (B/A)

(144000/288000) * 100

=50%

(291200/582400) * 100

=50%

(520960/1041920)*100

=50%

(3) Net Operating Income (Loss) :-

Net Operating Income (Loss) = Sals value – VC – Fixed cost

450 Units

910 Units

1628 Units

Sale value                                

(450 * $640)

=288000

(910 * $640)

=582400

(1628 * $640)

=1041920

(-) Variable Cost

144000

291200

520960

Contribution Margin             

144000

291200

520960

(-) Fixed Cost

198380

198380

198380

Net Operating Income (Loss)

(54380)

92820

322580